I realize some of these look easy but when you haven't had algebra for so long it looks hard to me/ I also appreciate your help, without you guys I would be Sc-----.
Please evaluate the following and if so can you explain your steps.
X^2 -5X + 7 when X = -2

Answer by ptaylor(1894)

X^2 -5X + 7 when X = -2
Step 1: (-2)^2=+4
Step 2: (-5)(-2)=+10
Step 3: Put it all back together:
+4+10+7=21
Hope this helps---ptaylor